While walking home, Mae casts a shadow of 1.15m when she is 0.75m away from a 4.42m high street lamp. How tall is Mae?
Answer by Alan3354(69443)

While walking home, Mae casts a shadow of 1.15m when she is 0.75m away from a 4.42m high street lamp. How tall is Mae?
--------
The distance from the pole to the far end of her shadow is 115 + 75 cm = 190 cm
---
h = her height in cms
442/190 = h/115